OpenStudy (shealah):

Find the circumcenter of triangle CAT, C(-10,10) A(-4,-2) and T(4,-2) algebraically.

OpenStudy (linknissan):

(-10/3,2)

OpenStudy (shealah):

I found the perpendicular bisector of CA is y=1/2x + 7.5 and the perpendicular bisector of CT is y=8/12x+6...

OpenStudy (linknissan):

no need here is ur prob solved ..cimcumcntr= (x1+x2+x3)/3 will giv u d X co ordinate of the circumcntr..similyar do it for y

OpenStudy (shealah):

i am supposed to find the perpendicular bisector for AT but it comes out 0 so i just ignore it???

OpenStudy (linknissan):

wer u told to do it algebriaclly or or solve with proofs??

OpenStudy (shealah):

algebraically.

OpenStudy (linknissan):

thn jus use the direct formula..